# Growth of Structural Lengthscale in Kob Andersen Binary Mixtures: Role of medium range order

### v1

@@ Abstract
A central and extensively debated question in glass physics concerns whether a single, growing lengthscale fundamentally controls glassy dynamics, particularly in systems lacking obvious structural motifs or medium range crystalline order (MRCO). In this work, we investigate structural and dynamical lengthscales in supercooled liquids using the Kob Andersen binary Lennard Jones (KALJ) model in two compositions: 80:20 and 60:40. We compute the dynamical lengthscale from displacement–displacement correlation functions and observe a consistent growth as temperature decreases. To explore the static counterpart, we use a structural order parameter (SOP) based on the mean field caging potential. While this SOP is known to predict short time dynamics effectively, its bare correlation function reveals minimal spatial growth. Motivated by recent findings that long time dynamics reflect collective rearrangements, we perform spatial coarse-graining of the SOP and identify an optimal lengthscale $L_{max}$ that maximises structure–dynamics correlation. We show that the structural correlation length derived from SOP coarse-grained over $L_{max}$ exhibits clear growth with cooling and closely tracks the dynamical lengthscale, especially for A particles in the 80:20 mixture and for both A and B particles in the 60:40 system. Our results reconcile the previously observed absence of static length growth in MRCO-free models like KALJ by highlighting the necessity of intermediate range structural descriptors. Furthermore, we find that the particles with larger structural length growth also correspond to species with latent crystallisation tendencies, suggesting a possible link between structural order, dynamics, and incipient crystallisation.
@@ I Introduction
Supercooled liquids exhibit a dramatic slowdown in dynamics upon cooling and ultimately form an amorphous solid, or glass Debenedetti2001 ; Cavagna2009 ; Berthier2011 ; Ediger2000 . Despite sustained progress, the microscopic mechanisms driving this glass transition remain debated. A central challenge is to rationalise the rapid increase in the dynamics, spanning many orders of magnitude, with a corresponding static structural signature tanakamrco_kawasaki_prl_2007 ; Coslovich2007 ; royall2008 ; tanakamrco_shintani_natphys_2008 ; Karmakar2009 ; Tanaka2010 ; keys2011 ; Coslovich2011 ; Sastry2014 ; bapst2020 ; Manoj_prl_2021 ; Sharma2022 . This issue has motivated extensive efforts to identify growing structural and dynamical correlation lengths, a key theme in modern glass physics shintani2006 ; tanakamrco_kawasaki_prl_2007 ; tanakamrco_critical_natmat_2010 ; Dunleavy2012 ; MalinsLifetimes2013 ; russo_tanaka ; Tarjus2016pre ; Tah_PRL_2018 .
Dynamical lengthscales are relatively well established, often quantified through multi-point correlation functions such as the four-point dynamic structure factor Karmakar2010 ; Berthier2011 ; Tah_dis_dis ; Paula2023 or the displacement–displacement correlation function  Donati1999PRL ; Tah_dis_dis ; Paula2023 . In contrast, static lengthscales remain elusive, largely because conventional pair correlation functions like radial distribution function ($g(r)$) exhibit little change upon supercooling.
@@
Over the past two decades, considerable effort has been devoted to identifying static lengthscales that grow alongside the well established dynamical correlation length in glass forming systems. Tanaka and coworkers have extensively studied the growth of static lengthscale in systems exhibiting medium range crystalline order (MRCO), such as those with Steinhardt bond orientational order parameterSteinhardt1983 in 3D and hexatic order in 2D and found that these systems show a stronger coupling between static and dynamic lengthscales tanakamrco_shintani_natphys_2008 ; tanakamrco_kawasaki_prl_2007 ; tanakamrco_kawasaki_JPCM_2010 ; tanakamrco_critical_natmat_2010 ; tanakamrco_leocmach_natcomm_2012 ; tanakamrco_boo_jncs_2012 ; tanakamrco_hu_pree_2016 . Note that although in 2D systems, the presence of MRCO is quite common and easy to detect, it is often tricky to find the same in 3D systems. Advanced techniques for measuring static lengthscales, such as the point-to-set (PTS) method, were designed to be order-agnostic biroli_thermodynamic_2008 ; Hocky2012 ; Biroli2013 ; Berthier2016 ; Tarjus2016pre . Tah et al. Tah_PRL_2018 investigated the PTS length in glassy systems, distinguishing between those with and without medium range crystalline order (MRCO). Their work revealed a sublinear scaling of the static PTS length with the dynamical correlation length for systems lacking MRCO, while for MRCO systems, these two lengths were found to grow commensurately. Royall and co-workers have shown that for polydisperse systems in deep supercooling, the static lengthscale connected to the defective icosahedra order grows modestly Hallett2018 .
@@
However, for MRCO-free systems like the 3D Kob Andersen binary Lennard Jones mixture, this correspondence often breaks down. Early studies by Marinari and Pitard Marinari2005 and La Nave et al. LaNave2006 linked cooperative relaxation to persistent clusters and inherent structure landscapes, respectively. Coslovich and Pastore used Voronoi tessellation to identify locally preferred structures (LFS) which are connected to the slow dynamics and showed that their cluster size grows with decreasing temperature Coslovich2007 . Royall and coworkers used their proposed Topological cluster classification (TCC) Malins2013JCP to identify LFS connected to slow particles and studied the growth of these clusters Malins2013JCP ; Dunleavy2012 ; MalinsLifetimes2013 ; Royall2014 ; Crowther2015 ; Hallett2018 ; royall2008 . Karmakar et al. KARMAKAR20121001 showed that the minimum eigenvalue of the hessian matrix $\lambda_{min}$ plays a crucial role in defining and understanding the static length at low temperature where an-harmonic effect is absent. Chakrabarty et al. Chakrabarty2017 introduced a new structure agnostic methodology called block analysis to study the static and dynamic lengthscales. Nevertheless, all these studies show that in systems without medium range crystalline order (MRCO) these structural lengthscales grow more slowly than their dynamical counterparts.
@@
In the KALJ model, the topological cluster classification (TCC) studies identified the 11A bicapped square-antiprism cluster, which is assumed to be a precursor of the CsCl structure, as an LFS associated with slow dynamics, but its spatial extent does not scale with the dynamical correlation length Malins2013JCP ; MalinsLifetimes2013 ; Royall2014 . More recent investigations suggest that in the KALJ model, the particles do not have a CsCl formation tendency between the two species, but the A-type particles show a tendency to form a face-centred cubic (fcc) structure Ujjwal_jcp_2016 . However, these motifs are not easily captured by standard bond-orientational or TCC-based descriptors. This disconnect raises an important question: does the lack of observed static lengthscale growth stem from a true absence of underlying order, or does it reflect the limitations of existing local structural metrics that fail to capture intermediate-range correlations?
@@
In this work, we revisit the Kob Andersen binary Lennard-Jones mixture, examining both the canonical 80:20 composition and a modified 60:40 variant. We begin by analysing the structural lengthscale using a recently developed structural order parameter (SOP) that has been shown to correlate well with particle dynamics Manoj_prl_2021 ; Sharma2022 . While this local bare SOP effectively captures short time dynamical behaviour, we find that it does not show a significant growth in spatial correlations upon supercooling and thus fails to reveal a growing static lengthscale. It is important to note that dynamical correlation lengths are typically extracted from long time, not short time, dynamics. Previous studies on both 2D and 3D polydisperse systems by Tong and Tanaka have demonstrated that for their generalised angular structural order, coarse-graining the order parameter in space provides better correlation with long time dynamics. They also showed that the optimal coarse-graining length, at which the structure-dynamics correlation is maximised, increases with decreasing temperature TongTanaka2018 ; TongTanaka2019 ; TongTanaka2020 ; Tanaka2025 . Study involving elastically collective nonlinear Langevin equation theory has shown that the inter relationship between alpha relaxation time and the local cage and collective elastic barriers and also other thermodynamic quantities, is present only when medium range order in the static length is taken into consideration in the theoryken_sch . The study further showed that the static length grows with density. Similarly, investigations using local volume MeiWang2022 , local caging potential mohit_wca , and order parameters derived via unsupervised machine learning QiuArun2025 have shown that meaningful correlations with long time dynamics emerge only after spatial coarse-graining of the static order parameter. Here, we demonstrate that the SOP, when coarse-grained over an optimal lengthscale, exhibits strong correlations with long time dynamics. Notably, the static lengthscale extracted from this coarse-grained SOP grows with decreasing temperature in a manner consistent with the dynamical lengthscale. This growth, for the first time is observed in the MRCO-free KALJ model, provides new evidence that medium range structural correlations, not purely local order, are essential to understanding the dynamical slowdown in glass-forming systems lacking conventional crystalline precursors.
@@
The organisation of the rest of the paper is as follows. Section II provides the simulation details. In Section III, we describe the methodology used in this work. Section IV presents the results. Section V presents a concluding discussion. Two appendices are included for additional details.
@@ II Simulation Details
where $\alpha,\beta\in\{\text{A},\text{B}\}$ denote the types of the interacting particles. The interaction parameters are set as follows: $\epsilon_{\mathrm{AA}}=1.0$, $\epsilon_{\mathrm{AB}}=1.5$, $\epsilon_{\mathrm{BB}}=0.5$, $\sigma_{\mathrm{AA}}=1.0$, $\sigma_{\mathrm{AB}}=0.8$, and $\sigma_{\mathrm{BB}}=0.88$. All quantities are expressed in reduced LJ units. The number density is fixed at $\rho=1.2$ for the 80:20 mixture and $\rho=1.41$ for the 60:40 mixture. Interactions are truncated and quadratically shifted at a cutoff distance of $r_{c}=2.5\sigma_{\mathrm{AB}}$ to ensure continuity of both the potential and its first derivative Stoddard1973 . Simulations are performed in the NVT ensemble using a Nosé-Hoover thermostat with a time step $\Delta t=0.005$. To minimize finite-size effects, we set the system size to $N=16{,}000$ for structural quantity calculations and $N=500{,}000$ for evaluating the four-point correlation function, following the approach outlined in Reference Karmakar2010 .
@@ III.1 Overlap function and four-point structure factor
The, four-point structure factor $S_{4}({\bf{q}},t)$, which characterises the spatial correlations of the dynamical fluctuations in the wavevector plane q is given by Karmakar2010
@@
where $\xi_{4}$ is the dynamic correlation length that grows on approaching the glass transition Karmakar2010 ; Paula2023 . For data analysis, we plot $1/S_{4}(q,\tau_{\alpha})$ versus $q^{2}$; from Eq.5 this should be linear, so an ordinary least-squares fit reliably extracts both the intercept $1/S_{4}(q\!\to\!0,\tau_{\alpha})$ and the slope $\xi_{4}^{2}/S_{4}(q\!\to\!0,\tau_{\alpha})$, which can be then used to obtain $\xi_{4}$.
@@ III.2 Displacement–displacement correlation function
To characterise the spatial extent of correlated motion in supercooled liquids, we use the displacement–displacement correlation function $g_{uu}(r,\tau_{\alpha})$, as studied earlier, Poole1998PhysicaA ; Donati1999PRL ; Paula2023 ; Hallett2018 .
It is defined as:
$g_{uu}(r,\tau_{\alpha})=\frac{\left\langle\sum_{\begin{subarray}{c}i,j=1\\ j\neq i\end{subarray}}^{N}u_{i}(\tau_{\alpha})u_{j}(\tau_{\alpha})\delta\left(r-|\mathbf{r}_{ij}(t)|\right)\right\rangle}{4\pi r^{2}\Delta rN\rho\langle u(\tau_{\alpha})\rangle^{2}}$ (6)
where $u_{i}(\tau_{\alpha})=|\mathbf{r}_{i}(t+\tau_{\alpha})-\mathbf{r}_{i}(t)|$ is the scalar displacement of particle $i$ over time $\tau_{\alpha}$.
To isolate the dynamic part from static density correlations, the computation of the excess correlation function is done by using
$\Gamma_{uu}(r,\tau_{\alpha}))=\frac{g_{uu}(r,\tau_{\alpha})}{g(r)}-1$ (7)
where g(r) is the radial distribution function and is defined as
$g(r)=\frac{\left\langle\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{N}\delta\left(r-|\mathbf{r}_{ij}(t)|\right)\right\rangle}{4\pi r^{2}\Delta r\rho N}$ (8)
where $\rho$ is the number density and N is the number of particles. An increasing $\Gamma_{uu}(r,\tau_{\alpha})$ with decreasing temperature indicates growing dynamical heterogeneity near the glass transition. Assuming that $\Gamma_{uu}(r,\tau_{\alpha})$ decays exponentially with distance, then the associated lengthscale can be extracted by integrating $\Gamma_{uu}(r,\tau_{\alpha})$ over space as shown in the equation belowPoole1998PhysicaA ,
[…]
yielding $\xi_{D}$, the characteristic dynamical correlation length.

### v2

@@ Abstract
A central and extensively debated question in glass physics concerns whether a single, growing lengthscale fundamentally controls glassy dynamics, particularly in systems lacking obvious structural motifs like the Kob Andersen binary Lennard Jones (KALJ) model. In this work, we investigate structural and dynamical lengthscales in supercooled liquids using KALJ model in two compositions: 80:20 and 60:40. We compute the dynamical lengthscale from displacement–displacement correlation functions and observe a consistent growth as temperature decreases. To explore the static counterpart, we use a structural order parameter (SOP) based on the mean field caging potential. While this SOP is known to predict short time dynamics effectively, its bare correlation function reveals minimal spatial growth. Motivated by recent findings that long time dynamics reflect collective rearrangements, we perform spatial coarse-graining of the SOP and identify an optimal lengthscale $L_{max}$ that maximises structure–dynamics correlation. We show that the structural correlation length derived from SOP coarse-grained over $L_{max}$ exhibits clear growth with cooling and closely tracks the dynamical lengthscale, especially for A particles in the 80:20 mixture and for both A and B particles in the 60:40 system. Our results reconcile the previously observed absence of static length growth in the KALJ model by highlighting the necessity of intermediate range structural descriptors. Furthermore, we find that the particles with larger structural length growth also correspond to species with latent crystallisation tendencies, suggesting a possible link between structural order, dynamics, and incipient crystallisation.
@@ I Introduction
Supercooled liquids exhibit a dramatic slowdown in dynamics upon cooling and ultimately form an amorphous solid, or glass Debenedetti2001 ; Cavagna2009 ; Berthier2011 ; Ediger2000 . Despite sustained progress, the microscopic mechanisms driving this glass transition remain debated. A central challenge is to rationalise the rapid increase in the dynamics, spanning many orders of magnitude, with a corresponding static structural signature tanakamrco_kawasaki_prl_2007 ; Coslovich2007 ; royall2008 ; tanakamrco_shintani_natphys_2008 ; Karmakar2009 ; tanakamrco_critical_natmat_2010 ; keys2011 ; Coslovich2011 ; Sastry2014 ; bapst2020 ; Manoj_prl_2021 ; Sharma2022 . This issue has motivated extensive efforts to identify growing structural and dynamical correlation lengths, a key theme in modern glass physics shintani2006 ; tanakamrco_kawasaki_prl_2007 ; tanakamrco_critical_natmat_2010 ; Dunleavy2012 ; Tarjus_simpleglass_2012 ; MalinsLifetimes2013 ; russo_tanaka ; Tah_PRL_2018 .
Dynamical lengthscales are relatively well established, often quantified through multi-point correlation functions such as the four-point dynamic structure factor Karmakar2010 ; Berthier2011 ; Tah_dis_dis ; Paula2023 or the displacement–displacement correlation function  Donati1999PRL ; Tah_dis_dis ; Paula2023 . Studies of dynamical lengthscale have also been extended in the deep supercooled region below the mode coupling theory (MCT) transition temperatureFlenner2013 ; Hallett2018 ; berthier_PRX_2022 . In some of these studies, it was shown that the growth of the dynamical lengthscale slows down in this deep supercooled regionFlenner2013 ; Hallett2018 .
@@
Over the past two decades, considerable effort has also been devoted to identifying static lengthscales that grow alongside the well established dynamical correlation length in glass forming systems. Tanaka and coworkers have extensively studied the growth of static lengthscale in systems exhibiting MRCO, such as those with Steinhardt bond orientational order parameterSteinhardt1983 in 3D and hexatic order in 2D and found that these systems show a stronger coupling between static and dynamic lengthscales tanakamrco_shintani_natphys_2008 ; tanakamrco_kawasaki_prl_2007 ; tanakamrco_kawasaki_JPCM_2010 ; tanakamrco_critical_natmat_2010 ; tanakamrco_leocmach_natcomm_2012 ; tanakamrco_boo_jncs_2012 ; tanakamrco_hu_pree_2016 .
Note that in 2D systems, the presence of MRCO, which is primarily a hexatic order, is quite common and easy to detect tanakamrco_critical_natmat_2010 ; russo_tanaka ; tanakamrco_kawasaki_prl_2007 ; TongTanaka2018 ; TarjusPRL2010 . However, in 3D systems, the MRCO is not so common and is obtained mostly in deep supercooling. In marginally polydisperse systems, these structures are face centered cubic (fcc) tanakamrco_leocmach_natcomm_2012 or hexagonal close packing tanakamrco_critical_natmat_2010 and in systems with higher polydispersity they are icosahedral or defective icosahedral structures Hallett2018 . Advanced techniques for measuring static lengthscales, such as the point-to-set (PTS) method, were designed to be order-agnostic biroli_thermodynamic_2008 ; Hocky2012 ; Biroli2013 ; Berthier2016 . However, Russo and Tanaka russo_tanaka showed that for a two dimensional polydisperse system exhibiting strong hexatic order, the PTS lengthscale fails to capture the underlying hexatic order that actually drives the slowing down of the dynamics. They suggested that the PTS length primarily reflects two-body correlations encoded in the radial distribution function. Tah et al. Tah_PRL_2018 investigated the PTS length, distinguishing between those with and without MRCO. Their work for 2D systems revealed a sublinear scaling of the static PTS length with the dynamical correlation length for systems lacking MRCO, while for MRCO systems, these two lengths were found to grow commensurately if the cavity was chosen in a way that the inside liquid mimics the thermodynamic properties of the bulk liquid. Royall and co-workers have shown that for polydisperse systems in deep supercooling, the static lengthscale connected to the locally favoured structure, which is a defective icosahedra order, grows almost in a similar fashion as the dynamical lengthscale Hallett2018 .
@@
However, for MRCO-free systems like the 3D Kob Andersen binary Lennard Jones mixture, this correspondence often breaks down. Early studies by Marinari and Pitard Marinari2005 and La Nave et al. LaNave2006 linked cooperative relaxation to persistent clusters and inherent structure landscapes, respectively. Coslovich and Pastore used Voronoi tessellation to identify locally preferred structures (LPS), which are connected to the slow dynamics, and showed that their cluster size grows with decreasing temperature Coslovich2007 . Royall and coworkers used their proposed Topological cluster classification (TCC) Malins2013JCP to identify LFS connected to slow particles and studied the growth of these clusters Malins2013JCP ; Dunleavy2012 ; MalinsLifetimes2013 ; Royall2014 ; Crowther2015 ; Hallett2018 ; royall2008 . Karmakar et al. KARMAKAR20121001 showed that the minimum eigenvalue of the hessian matrix $\lambda_{min}$ plays a crucial role in defining and understanding the static length at low temperature where an-harmonic effect is absent. Chakrabarty et al. Chakrabarty2017 introduced a new structure agnostic methodology called block analysis to study the static and dynamic lengthscales. The PTS correlation length has also been obtained for many systems Hocky2012 ; Berthier2016 ; russo_tanaka . Nevertheless, all these studies show that in systems without MRCO, these structural lengthscales grow more slowly than their dynamical counterparts. Even studies exploring deep supercooled regime below the MCT transition temperature, where the activated dynamics is relevant, have shown that the dynamical lengthscale grows faster than the static PTS lengthscale berthier_PRX_2022 .
@@
In the KALJ binary 80:20 mixture of A and B particles, Fernandez and Horowell have shown that although the lattice energy of the total system is lowest when the system crystalizes into CsCl structure between the A and the B particles and the rest of the A particles form face-centered cubic (fcc) structure, isolated CsCl structure was found to be less stable compared to isolated square antiprism structure with B particles at the center FernandezHarrowell_PRE_2003 . The latter structure is related to the $Al_{2}Cu$ crystal. The presence of this structure was observed by Crowther et al. using their TCC method Malins2013JCP and referred to as 11A structure Crowther2015 . This 11A structure was found to be an LFS associated with slow dynamics MalinsLifetimes2013 ; Royall2014 , but due to compositional frustration, its spatial extent was found not to grow Crowther2015 , which led to the decoupling of the growth of the static and the dynamical lengthscale MalinsLifetimes2013 ; Royall2014 . More recent investigations suggest that in the KALJ model, the particles do not have a CsCl formation tendency between the two species, but all the A-type particles show a tendency to form a distorted fcc structure where the distortion comes due to the presence of the B particles in between Ujjwal_jcp_2016 . However, these motifs are not easily captured by standard bond-orientational order or TCC-based descriptors. This disconnect raises an important question: Does the lack of observed static lengthscale growth stem from a true absence of underlying order, or does it reflect the limitations of existing local structural metrics?
@@
In this work, we revisit the Kob Andersen binary Lennard-Jones mixture, examining both the canonical 80:20 composition and a modified 60:40 variant. We begin by analysing the structural lengthscale using a recently developed structural order parameter (SOP) that has been shown to correlate well with particle dynamics Manoj_prl_2021 ; Sharma2022 . However, we find that it does not show a significant growth in spatial correlations upon supercooling and thus fails to reveal a growing static lengthscale. It is important to note that dynamical correlation lengths are typically extracted from long time and not short time, dynamics. Whereas for most order parameters, the success is decided by correlating it with the short time dynamics liu_nature ; Sharma2022 ; Zhang2020 . Recently, Tong and Tanaka have proposed that for systems with and without MRCO, when the relevant order parameter is coarse grained in space, it provides better correlation with the long time dynamics TongTanaka2018 . According to their study, the optimal coarse-graining length, at which the structure-dynamics correlation is maximised, increases with decreasing temperature, and this growth is similar to the growth of the dynamical lengthscale  TongTanaka2018 ; TongTanaka2019 ; TongTanaka2020 ; Tanaka2025 . There are also other studies which highlight the role of medium range static order in describing the dynamics. A study involving elastically collective nonlinear Langevin equation theory has shown that medium-range order is important in understanding structure dynamics correlation, and the static length then grows with density ken_sch . Similarly, investigations using local volume MeiWang2022 , local caging potential mohit_wca , and order parameters derived via unsupervised machine learning QiuArun2025 have shown that meaningful correlations with long time dynamics emerge only after spatial coarse-graining of the static order parameter. Here, we demonstrate that the SOP, when coarse-grained over an optimal lengthscale, exhibits strong correlations with long time dynamics. Notably, the static lengthscale extracted from this coarse-grained SOP grows with decreasing temperature in a manner consistent with the dynamical lengthscale. While medium range structural correlations in MRCO free models have previously been detected in additive binary mixtures TongTanaka2018 ; TongTanaka2019 ; TongTanaka2020 ; Tanaka2025 , our study successfully identifies such correlations in one of the most standard glass forming models, the non-additive Kob Andersen model. The result emphasizes the previous findings TongTanaka2018 ; TongTanaka2019 ; TongTanaka2020 ; Tanaka2025 ; ken_sch ; MeiWang2022 ; QiuArun2025 that medium range structural correlations and not purely local order, are essential in understanding the dynamical slowdown in glass-forming systems.
@@
The organisation of the rest of the paper is as follows. Section II provides the simulation details. In Section III, we describe the methodology used in this work. Section IV presents the results. Section V presents a concluding discussion. Five appendices are included for additional details.
@@ II Simulation Details
where $\alpha,\beta\in\{\text{A},\text{B}\}$ denote the types of the interacting particles. The interaction parameters are set as follows: $\epsilon_{\mathrm{AA}}=1.0$, $\epsilon_{\mathrm{AB}}=1.5$, $\epsilon_{\mathrm{BB}}=0.5$, $\sigma_{\mathrm{AA}}=1.0$, $\sigma_{\mathrm{AB}}=0.8$, and $\sigma_{\mathrm{BB}}=0.88$. All quantities are expressed in reduced LJ units. The number density is fixed at $\rho=1.2$ for the 80:20 mixture and $\rho=1.41$ for the 60:40 mixture. Interactions are truncated and quadratically shifted at a cutoff distance of $r_{c}=2.5\sigma_{\mathrm{AB}}$ to ensure continuity of both the potential and its first derivative Stoddard1973 . Simulations are performed in the NVT ensemble using a Nosé-Hoover thermostat with a time step $\Delta t=0.005$. To minimize finite-size effects, we set the system size to N=16, [^0] N=16{,}000 for structural quantity calculations and N=500, [^0] N=500{,}000 for evaluating the four-point correlation function, following the approach outlined in Reference Karmakar2010 .
@@ III.1 Overlap function and four-point structure factor
The four-point structure factor $S_{4}({\bf{q}},t)$, which characterises the spatial correlations of the dynamical fluctuations in the wavevector plane q, is given by Karmakar2010
@@
where $\xi_{4}$ is the dynamic correlation length that grows on approaching the glass transition Karmakar2010 ; Paula2023 . For data analysis, we plot $1/S_{4}(q,\tau_{\alpha})$ versus $q^{2}$; from Eq.5 this should be linear, so an ordinary least-squares fit reliably extracts both the intercept $1/S_{4}(q\!\to\!0,\tau_{\alpha})$ and the slope $\xi_{4}^{2}/S_{4}(q\!\to\!0,\tau_{\alpha})$, which can be then used to obtain $\xi_{4}$. For this analysis, we use the 500,000 particle system, which allows us to explore smaller q ranges. However, in this small q range, for better averaging, we use a long run length of 200,000 time steps. We fit the data in the range q=0.10.6q=0.1-0.6.
@@ III.2 Displacement–displacement correlation function
To characterise the spatial extent of correlated motion in supercooled liquids, we use the excess displacement–displacement correlation function $g_{uu}(r,\tau_{\alpha})=\bigl\langle u(0)\,u(r)\bigr\rangle$ which is defined as, Poole1998PhysicaA ; Donati1999PRL ; Paula2023 ; Hallett2018 .
$\bigl\langle u(0)\,u(r)\bigr\rangle=\frac{\Bigl\langle\displaystyle\sum_{i,j}u(\tau_{\alpha})_{i}\,u(\tau_{\alpha})_{j}\,\delta\!\bigl(r-\lvert\mathbf{r}_{ij}\rvert\bigr)\Bigr\rangle}{4\pi r^{2}\,\Delta r\,\rho\,Ng(r)}$ (6)
where $u_{i}(\tau_{\alpha})=|\mathbf{r}_{i}(t+\tau_{\alpha})-\mathbf{r}_{i}(t)|$ is the scalar displacement of particle $i$ over time $\tau_{\alpha}$. g(r) is the radial distribution function (rdf) and is defined as
$g(r)=\frac{\left\langle\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{N}\delta\left(r-|\mathbf{r}_{ij}|\right)\right\rangle}{4\pi r^{2}\Delta r\rho N}$ (7)
where $\rho$ is the number density and N is the number of particles. The normalised excess displacement-displacement correlation function is given by LandauSchuttler1999
$\Gamma_{uu}(r,\tau_{\alpha})=\frac{\bigl\langle u(0)\,u(r)\bigr\rangle-\bigl\langle u(\tau_{\alpha})\bigr\rangle^{2}}{\bigl\langle\bigl(u(\tau_{\alpha})\bigr)^{2}\bigr\rangle-\bigl\langle u(\tau_{\alpha})\bigr\rangle^{2}}$ (8)
An increasing $\Gamma_{uu}(r,\tau_{\alpha})$ with decreasing temperature indicates growing dynamical heterogeneity near the glass transition. Assuming that $\Gamma_{uu}(r,\tau_{\alpha})$ decays exponentially with distance, then the associated lengthscale can be extracted by integrating $\Gamma_{uu}(r,\tau_{\alpha})$ over space as shown in the equation belowPoole1998PhysicaA ,
$\xi_{D}=\int_{0}^{\infty}\Gamma_{uu}(r,\tau_{\alpha})\,dr$ (9)
yielding $\xi_{D}$, the characteristic dynamical correlation length. For this calculation, we take a 16,000 particle system. The averaging is done over a run length of 950,000 timesteps for 4 configurations.